• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Annand

    Log entry time 07:59:36 on February 5,2002

    Entry number 81851

    keyword=Owl Shift Summary

    Owl Shift summary
    John Annand (SL)
    Igor Rachek (TO)
    Michael Roedelbronn (3)
    Vincent Sukovsky (4)

    RCS Kinematics 5C(e) Cal 26 deg, 14.5m, HRSL 32.66 deg, 2.772 GeV/c
    LH2 RAD, RCS Magnet ON (600A), Veto OFF, 15uA continued

    00.00 2089 HV problem calorimeter starting 12.08 approx
    00.35 2090 HV problem calorimeter, CODA crash (no end summary)
    01.10 Request controlled access to cycle power to TS2 and ROC4.
    01.42 2091...more CODA problems
    TS1 reset from control room....finally CODA seems to run again.
    01.55 2092 Hopefully restart 5C(e) production in earnest
    02.22 2093 Continue 5C(e)
    02.51 2094 Continue 5C(e)
    03.15 2095 Continue 5C(e)
    03.40 2096 Continue 5C(e)
    04.06 2097 Continue 5C(e)
    04.28 Short break for QE measurement by MCC
    04.41 2098 Continue 5C(e)
    05.06 2099 Continue 5C(e) Accumulated charge 0.4C approx at end of run

    05.28 Start 5C(f) RCS Magnet OFF
    05.28 2100 Junk TS2 crash...reset, restart CODA
    05.32 2101 Start 5C(f) production
    05.54 2102 Continue 5C(f)
    06.23 2102 Continue 5C(f) Accumulated 0.04C approx at end of run

    06.37 Start 5C(g) OPTICS target 25uA, Raster OFF
    06.37 2103 Start 5C(g) production....forgot to SAVE Prescale 3 = 1
    06.55 2104 5C(g) production Ps3 = 1
    07.07 Beam OFF